# Reinforcement Learning Teachers of Test Time Scaling

### v1

@@ B.4 Student evaluation
As described in Section 4.1, our main evaluation consists of three graduate and competition-level tasks on math and natural science domains. In particular, these include AIME24 [of America, 2024], the set of problems used for the American Invitational Mathematics Examination; MATH 500 [Hendrycks et al., 2021], the set of problems selected by [Lightman et al., 2023] from the canonical competition math benchmark; and GPQA Diamond [Rein et al., 2024], the set of diamond difficulty problems on natural science topics from the Graduate-level Google-proof Q&A benchmark. Additionally, in Appendix C, we also extend our set of experiments to include additional challenging coding and multilingual domains. In particular, we consider LiveCodeBench [Jain et al., 2024], a contamination-free set of coding challenge problems continuously collected from several prominent online hosting platforms; and OlympiadBench [He et al., 2024], a set of olympiad-level bilingual problems in English and Chinese from past math and physics competitions.
We evaluate on all the above benchmarks using Lighteval [Fourrier et al., 2023], a library available under the MIT license. In all our results, we report the completion accuracy of each of our students for a single generated completion, as also reported in prior work [Muennighoff et al., 2025, Li et al., 2025a, Labs, 2025]. Furthermore, we also ensure consistency by re-using the task implementation code, including the system prompt, provided by our baselines [Team, 2025b], which is available under an Apache 2.0 License. For the same reason, we do not modify any of the existing evaluation generation hyperparameters from the suggested settings used in their evaluation, which we report in Table 7.
@@ C.1 Coding and multilingual reasoning
We extend our main set of experiments from Section 4, focusing on graduate and competition-level tasks on math and natural science domains, comparing the effectiveness of the reasoning traces from our 7B parameter RLT with traditional distillation pipelines. We consider challenging coding and multilingual domains, which are less aligned with the training and distillation set of questions employed. In particular, these new benchmarks include LiveCodeBench (LCB) [Jain et al., 2024], a contamination-free set of coding challenge problems continuously collected from several prominent online hosting platforms across three difficulty categories; and OlympiadBench [He et al., 2024], a set of olympiad-level bilingual problems in English and Chinese from past math and physics competitions.
In Table 8, we provide a comparison of reported results from our baselines [Li et al., 2025a, Labs, 2025] and our fine-tuned student models using our 7B RLT traces. We also re-collected the performance on OlympiadBench of the Qwen2.5-7B-Instruct model and the Bespoke-7B baseline fine-tuned from the postprocessed R1 traces, as they were omitted in the reference results from prior work. For LiveCodeBench, we report both the performance on the “hard” difficulty set of problems and the average weighted performance. This was obtained by weighting the performances of the models in the “easy,” “medium,” and “hard” sets of LiveCodeBench problems, by the relative number of problems in each.
@@ Appendix D Extended analysis
Figure 9: Full formatted distillation data sample using a postprocessed DeepSeek R1 reasoning trace from the Bespoke dataset [Labs, 2025].
@@ D.1 RLT reward and R1 traces analysis
For these experiments, we train entirely new 7B RLTs following the hyperparameters from Table 3 but setting $\lambda=0$ (see Equation 5) or $\alpha=0$ (see Equations 4 and 3), respectively for the thought tokens KL reward and min/max reduction ablations. We then construct student distillation datasets, using the same full set of starting question-solution pairs as our recent state-of-the-art baselines [Li et al., 2025a, Labs, 2025] as considered in our main Section 4 experiments. We compare the generated datasets and students after distillation with these ablated teachers to our original 7B RLT trained with the full RLT reward and our strongest distillation baseline using postprocessed R1 reasoning traces [Li et al., 2025a, Labs, 2025].
First, we focus on the effects of these ablations in terms of the traces’ content as compared to our original RLT’s traces and the postprocessed R1 pipeline. As shown in Figure 8, the length of the reasoning traces is greatly affected by our ablations. When looking at our original RLT with full rewards, the average length of the produced reasoning trace is 39% higher than the original curated R1 traces in the Bespoke dataset [Labs, 2025]. This is consistent with our analysis from Section 4.5 and further highlighted by the full examples shown in Figures 9 and 10, showing how our 7B RLT often includes alternative verification steps and approaches not considered by traditional pipelines that do not optimize directly for downstream distillation. Furthermore, ablating each term in our reward leads to the concrete unwarranted effects described in Section 3.3, allowing the RL optimization procedure to find “shortcuts” to maximize $r^{\mathrm{RLT}}$ that hurt the quality of the reasoning traces. In particular, without the thought tokens KL reward $r^{\mathrm{KL}}$, our RLT cannot differentiate between explanations that guide the student step-by-step and those that increase the solution’s likelihood without a logical path that can be learned from. Thus, as shown in Figure 11, this ablation leads to a teacher that only learns to repeat the solution tokens themselves in its explanation to exploit the repetition tendency of pretrained student LMs, with the average length of its output reasoning traces dramatically dropping. Moreover, in our second ablation of the min/max reward reduction terms, the rewards become effectively biased by the length of the reasoning trace, leading the teacher to prefer long explanations only to reduce the influence on $r^{\mathrm{KL}}$ of hard but necessary individual logical steps. As a consequence of this bias, the average number of think tokens in the teacher’s reasoning traces almost doubles from the postprocessed R1 traces and, as shown in Figure 12, their content starts including many additional unnecessary steps that are just semantical repetitions of each other as learning progresses.
@@ E.1 Limitations and unexplored directions
This work’s purpose was to introduce a new class of Reinforcement-Learned Teachers designed to avoid the exploration challenge of sparse rewards and align the optimization of RL-trained LMs with the test-time goal of downstream distillation. However, there are still several limitations and improvements which we hope will be tackled in future extensions. First, to output explanations, the RLT framework relies on access to the ground-truth solutions. Hence, when used with datasets and domains where this information might not be available or not be practical to recover without querying LMs, small RLTs might have to again rely on larger models, even though still to a lesser extent than prior distillation pipelines. Similarly, as described in Section 4.1 and Appendix A, our current training recipe does not exclusively involve RL, with an initial phase to familiarize our small model with its new teaching format and role, again relying on some level of initial access to pre-collected reasoning examples that can be used accordingly. We also note that RLT training makes use of an additional student model for computing the rewards. In practice, however, we note the downsides from this were minimal, as simple parameter offloading removed any potential memory burden of having this model on GPU memory during backpropagation, and actual training time was dominated ($90\%$) by the costs of long-context autoregressive generation with our computational setup.

### v2

@@ B.4 Student evaluation
As described in Section 4.1, our main evaluation consists of three graduate and competition-level tasks on math and natural science domains. In particular, these include AIME24 [23], the set of problems used for the American Invitational Mathematics Examination; MATH 500 [24], the set of problems selected by [25] from the canonical competition math benchmark; and GPQA Diamond [26], the set of diamond difficulty problems on natural science topics from the Graduate-level Google-proof Q&A benchmark. Additionally, in Appendix C, we also extend our set of experiments to include additional challenging coding and multilingual domains. In particular, we consider LiveCodeBench [46], a contamination-free set of coding challenge problems continuously collected from several prominent online hosting platforms; and OlympiadBench [47], a set of olympiad-level bilingual problems in English and Chinese from past math and physics competitions.
We evaluate on all the above benchmarks using Lighteval [27], a library available under the MIT license. In all our results, we report the completion accuracy of each of our students for a single generated completion, as also reported in prior work [8, 11, 12]. Furthermore, we also ensure consistency by re-using the task implementation code, including the system prompt, provided by our baselines [31], which is available under an Apache 2.0 License. For the same reason, we do not modify any of the existing evaluation generation hyperparameters from the suggested settings used in their evaluation, which we report in Table 7.
@@ C.1 Coding and multilingual reasoning
We extend our main set of experiments from Section 4, focusing on graduate and competition-level tasks on math and natural science domains, comparing the effectiveness of the reasoning traces from our 7B parameter RLT with traditional distillation pipelines. We consider challenging coding and multilingual domains, which are less aligned with the training and distillation set of questions employed. In particular, these new benchmarks include LiveCodeBench (LCB) [46], a contamination-free set of coding challenge problems continuously collected from several prominent online hosting platforms across three difficulty categories; and OlympiadBench [47], a set of olympiad-level bilingual problems in English and Chinese from past math and physics competitions.
In Table 8, we provide a comparison of reported results from our baselines [11, 12] and our fine-tuned student models using our 7B RLT traces. We also re-collected the performance on OlympiadBench of the Qwen2.5-7B-Instruct model and the Bespoke-7B baseline fine-tuned from the postprocessed R1 traces, as they were omitted in the reference results from prior work. For LiveCodeBench, we report both the performance on the “hard” difficulty set of problems and the average weighted performance. This was obtained by weighting the performances of the models in the “easy,” “medium,” and “hard” sets of LiveCodeBench problems, by the relative number of problems in each.
@@ Appendix D Extended analysis
Figure 9: Full formatted distillation data sample using a postprocessed DeepSeek R1 reasoning trace from the Bespoke dataset [12].
@@ D.1 RLT reward and R1 traces analysis
For these experiments, we train entirely new 7B RLTs following the hyperparameters from Table 3 but setting $\lambda=0$ (see Equation 5) or $\alpha=0$ (see Equations 4 and 3), respectively for the thought tokens KL reward and min/max reduction ablations. We then construct student distillation datasets, using the same full set of starting question-solution pairs as our recent state-of-the-art baselines [11, 12] as considered in our main Section 4 experiments. We compare the generated datasets and students after distillation with these ablated teachers to our original 7B RLT trained with the full RLT reward and our strongest distillation baseline using postprocessed R1 reasoning traces [11, 12].
First, we focus on the effects of these ablations in terms of the traces’ content as compared to our original RLT’s traces and the postprocessed R1 pipeline. As shown in Figure 8, the length of the reasoning traces is greatly affected by our ablations. When looking at our original RLT with full rewards, the average length of the produced reasoning trace is 39% higher than the original curated R1 traces in the Bespoke dataset [12]. This is consistent with our analysis from Section 4.5 and further highlighted by the full examples shown in Figures 9 and 10, showing how our 7B RLT often includes alternative verification steps and approaches not considered by traditional pipelines that do not optimize directly for downstream distillation. Furthermore, ablating each term in our reward leads to the concrete unwarranted effects described in Section 3.3, allowing the RL optimization procedure to find “shortcuts” to maximize $r^{\mathrm{RLT}}$ that hurt the quality of the reasoning traces. In particular, without the thought tokens KL reward $r^{\mathrm{KL}}$, our RLT cannot differentiate between explanations that guide the student step-by-step and those that increase the solution’s likelihood without a logical path that can be learned from. Thus, as shown in Figure 11, this ablation leads to a teacher that only learns to repeat the solution tokens themselves in its explanation to exploit the repetition tendency of pretrained student LMs, with the average length of its output reasoning traces dramatically dropping. Moreover, in our second ablation of the min/max reward reduction terms, the rewards become effectively biased by the length of the reasoning trace, leading the teacher to prefer long explanations only to reduce the influence on $r^{\mathrm{KL}}$ of hard but necessary individual logical steps. As a consequence of this bias, the average number of think tokens in the teacher’s reasoning traces almost doubles from the postprocessed R1 traces and, as shown in Figure 12, their content starts including many additional unnecessary steps that are just semantical repetitions of each other as learning progresses.
@@ E.1 Limitations and unexplored directions
This work’s purpose was to introduce a new class of Reinforcement-Learned Teachers designed to avoid the exploration challenge of sparse rewards and align the optimization of RL-trained LMs with the test-time goal of downstream distillation. However, there are still several limitations and improvements which we hope will be tackled in future extensions. First, to output explanations, the RLT framework relies on access to the ground-truth solutions. Hence, when used with datasets and domains where this information might not be available or not be practical to recover without querying LMs, small RLTs might have to again rely on larger models, even though still to a lesser extent than prior distillation pipelines. Similarly, as described in Section 4.1 and Appendix A, our current training recipe does not exclusively involve RL, with an initial phase to familiarize our small model with its new teaching format and role, again relying on some level of initial access to pre-collected reasoning examples that can be used accordingly. We also note that RLT training makes use of an additional student model for computing the rewards. In practice, however, we note the downsides from this were minimal, as simple parameter offloading removed any potential memory burden of having this model on GPU memory during backpropagation, and actual training time was dominated ($>90\%$) by the costs of long-context autoregressive generation with our computational setup.